Find the cube root of each of the following numbers by prime factorisation method.

 

(vii) 110592

(viii) 46656

 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(viii) 46656

Solution:

46656 = 2×2×2×2×2×2×3×3×3×3×3×3

By grouping the factors in triplets of equal factors,

46656 = (2×2×2)×(2×2×2)×(3×3×3)×(3×3×3)

Here, 46656 can be grouped into triplets of equal factors,

∴ 46656 = (2×2×3×3) = 36

Hence, 36 is cube root of 46656.



Thanks